OpenStudy (anonymous):

The sides of a triangle have side lengths equal to a,b and√ (a^2−ab+b^2). What is the measure of the angle (in degrees) between the longest and shortest side?

OpenStudy (anonymous):

a>b or b>a? This is quite important?

OpenStudy (anonymous):

you can assume

OpenStudy (anonymous):

If a>b then, b<√ (a^2−ab+b^2) <a If b>a then, a<√ (a^2−ab+b^2) <b Now, let the angle be x then, ( √ (a^2−ab+b^2))^2 = a^2+b2-2ab cosx a^2 -ab +b^2=a^2+b^2-2ab cosx cosx=1/2 x=60